Rita hikes 3 and a half miles on a nature trail. How many half mile long sections of the trail did Rita hike?
Answer by stanbon(75887)

Rita hikes 3 and a half miles on a nature trail. How many half mile long sections of the trail did Rita hike?
-----------------------------
Answer:: 3/(1/2) =3(2/1) = 6
===============
Cheers,
Stan H.
===============